ANALYTICAL DESCRIPTION OF FREQUENCIES, DIPOLE STRENGTHS AND ROTATIONAL STRENGTHS FOR FUNDAMENTAL AND OVERTONE CH-STRETCHING TRANSITIONS

G. LONGHI, R. GANGEMI, AND \underline{S. ABBATE}, Dipartimento di Scienze Biomediche e Biotecnologie, Universita di Brescia, via Valsabbina 19, 25123 Brescia, Italy.

The normal mode to local mode transition has been described in the past by basic arguments centered on symmetry and on perturbative-type numerical quantum mechanical calculations ^{a}^{b}. In this work we investigate the absorption and Vibrational Circular Dichroism (VCD) spectra for a two degrees of freedom model of an HCCH chiral fragment endowed with C_{2}-symmetry, for the fundamental (Δv=1) and first two overtone regions Δv=2,3). We include electrical anharmonicity ^{c} in addition to mechanical anharmonicity, and deal with them in the framework of the Van Vleck contact transformation theory ^{d}^{e}. By making use of an algebraic manipulator (Maple) we are able to derive useful analytical expressions for frequencies, dipole strengths and rotational strengths for Δv=1,2,3.

aK.K. Lehmann, J.Chem.Phys. 79, 1098 (1983)

bI.M. Mills, A.G. Robiette, Mol. Phys. 56, 743 (1985)

cG. Longhi, L. Ricard, S. Abbate, G. Zerbi, J. Chem. Phys., 88, 6733 (1988)

dG. Amat, H.H. Nielsen, G. Tarrago, Rotation-Vibration in Polyatomic Molecules, M. Dekker, NY,(1971)

eX.G. Wang, E.L. Sibert III, J. Chem. Phys., 113, 5384 (2000)